QUESTION 1 If modular n = 74, e = 5, what is the multiplicative inverse of e?
Answer #1

e=5, Po=0, P,=1 n = 74 1 122 e-1 mod 74 El. Р i P+P; -, •q mod in 1-2 stepo: 74 = 14 (5) + 4 ; Poco 5=1 (4) + P=1 Step 1: ste